OpenStudy (anonymous):

Please solve this 4 the grade algebra problem? A group of students calculated their average score at a spelling bee. They realized that if one of them scored 9 more points , their average score will be 81 points. if one of them scored 3 points less, their average score will be 78 points. How many students were in the group?

OpenStudy (anonymous):

Let x = the number of students 81x - 9 = 78x + 3 Logic : number of students multiplied by the average will give you total points. You're given the changes (9 more than avg and 3 less than avg). So you subtract 9 from the first avg to get the unchanged avg and add 3 to the second avg to get the unchanged avg. The unchanged avgs should be the same so you set them equal to one another. Solving for x 81x - 9 - 78x = 78x + 3 - 78x 3x = 12 x = 4